# A Recipe For Obtaining Algebraic Addition Theorems Of The Weierstrass Elliptic Function

### v1

@@ 3 Applications And Discussion
Hence subject to the condition that $z,3z,4z\notin\Lambda_{\Omega}$, we have the following triplication formula,
$\wp(3z)=\frac{1}{9}\left(\frac{\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)-\wp^{\prime}(z)\wp^{(4)}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}\right)^{2}+\frac{1}{3}\frac{\wp^{\prime\prime}(z)\wp^{(4)}(z)-\wp^{\prime\prime\prime}(z)^{2}}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}-3\wp(z).$ (10)
If we assume that in addition $\wp(z)\neq 0$, we get the more complicated ones
@@
×$\times\left(g_{2}+2\frac{\wp(z)\wp^{\prime\prime\prime}(z)^{2}+\wp^{\prime}(z)^{2}\wp^{(4)}(z)+\wp^{\prime\prime}(z)^{3}-\wp(z)\wp^{\prime\prime}(z)\wp^{(4)}(z)-2\wp^{\prime}(z)\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}\right)$
@@
We obviously recommend the formula (10) for practical use. Above equations can be written only in terms of $\wp(z)$ utilizing the following
′′′$\wp^{\prime\prime\prime}=12\wp^{\prime}\wp\qquad\text{and}\qquad\wp^{(4)}=120\wp^{3}-18g_{2}\wp-12g_{3}.$
@@
where $\wp=\wp(z)$. For example equation (10) is equivalent to
@@
We remark that one might obtain simpler formulas such as (8) and (9) for $\wp(z_{1}+\cdots+z_{\ell})$ by considering the system
@@
This corresponds to the case $n_{1}=\ell-1$, $\gamma_{1}=1$, $(k_{1},\cdots,k_{\ell-1},k_{\ell})=(\ell-2,\cdots,0,-2)$. Similarly, the determinant in condition (3) can be calculated by Frobenius-Stickelberger formula and is equivalent to the statement that $z_{i}$ are distinct $\bmod\Lambda_{\Omega}$ with $z\notin\Lambda$. Taking the limits $z_{1},\cdots,z_{\ell}\to s$, we get the $\ell$-multiplication formula for $\wp(s)$.

### v2

@@ 3 Applications and Discussion
Hence we have the following triplication formula,
$\wp(3z)=\frac{1}{9}\left(\frac{\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)-\wp^{\prime}(z)\wp^{(4)}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}\right)^{2}+\frac{1}{3}\frac{\wp^{\prime\prime}(z)\wp^{(4)}(z)-\wp^{\prime\prime\prime}(z)^{2}}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}-3\wp(z).$ (3.3)
We furthermore get the more complicated ones
@@
$\times\left(g_{2}+2\frac{\wp(z)\wp^{\prime\prime\prime}(z)^{2}+\wp^{\prime}(z)^{2}\wp^{(4)}(z)+\wp^{\prime\prime}(z)^{3}-\wp(z)\wp^{\prime\prime}(z)\wp^{(4)}(z)-2\wp^{\prime}(z)\wp^{\prime\prime}(z)\wp^{\prime\prime\prime}(z)}{\wp^{\prime\prime}(z)^{2}-\wp^{\prime}(z)\wp^{\prime\prime\prime}(z)}\right)$
@@
We obviously recommend the formula (3.3) for practical use. Above equations can be written only in terms of $\wp(z)$ utilizing the following
$\wp^{\prime\prime\prime}=12\wp^{\prime}\wp\qquad\text{and}\qquad\wp^{(4)}=120\wp^{3}-18g_{2}\wp-12g_{3}.$
@@
where $\wp=\wp(z)$. For example equation (3.3) is equivalent to
@@
We remark that one might obtain simpler formulas such as (3.1) and (3.2) for $\wp(z_{1}+\cdots+z_{\ell})$ by considering the system
@@
This corresponds to the case $n_{1}=\ell-1$, $\gamma_{1}=1$, $(k_{1},\cdots,k_{\ell-1},k_{\ell})=(\ell-2,\cdots,0,-2)$. Similarly, the determinant in condition (2.1) can be calculated by Frobenius-Stickelberger formula and is equivalent to the statement that $z_{i}$ are distinct $\bmod\Lambda_{\Omega}$ with $z\notin\Lambda$. Taking the limits $z_{1},\cdots,z_{\ell}\to s$, we get the $\ell$-multiplication formula for $\wp(s)$.
Disclosure Statement. The author(s) report no potential conflict of interest.